Monica bought a $2000 4K Ultra HD TV for 15% down, with the balance to be paid with interest at 14% compounded monthly in six equal monthly payments. Construct the full amortization schedule for the debt. Calculate the total interest paid. (Do not round intermediate calculations. Round your answers to 2 decimal places. Leave no cells blank - be certain to enter "0" wherever required.)
Answer #1

Solution:

Loan Amount=$2000*(1-0.15)=$1700

Monthly interest rate(r)=0.14/12=0.012

no. of payment(n)=6

Monthly payment=Loan Amount[r(1+r)^n/(1+r)^n-1]

=$1700[0.012(1+0.012)^6/(1+0.012)^6-1]

=$295

Amortzation schedule

No. of payment Opening balance Monthly payment Interest Principal Closing balance
1 $1700 $295 $19.83 275.17 $1424.83
2 $1424.83 $295 $16.62 $278.38 $1146.45
3 $1146.45 $295 $13.38 $281.62 $864.83
4 $864.83 $295 $10.09 $284.91 $579.92
5 $579.92 $295 $6.77 $288.23 $291.65
6 $291.65 $295 $3.4 $291.65 0

Total Interest paid=($295*6)-$1700

=$70